# Enhanced quantum metric due to vacancies in graphene

### v1

@@ Abstract
The two-dimensionality and linear Dirac dispersion of graphene generate remarkable physical properties Novoselov et al. (2004, 2005); Castro Neto et al. (2009); Wehling et al. (2014). However, the vanishing density of states at half-filling, at the Dirac point, causes a natural insensitivity to electron interactions and its generated many-body instabilities, otherwise key concepts in condensed matter. But, it makes graphene sensitive to the presence of structural defects Castro Neto et al. (2009); Banhart et al. (2011); Gao et al. (2018). Defects often create localized states that affects the conductivity, including in graphene Stauber et al. (2007); Wehling et al. (2010), and can even turn a metal into an insulator Evers and Mirlin (2008), while defect states that are not fully localized have been shown to favor some long-range correlations Feigel’man et al. (2007); Burmistrov et al. (2012, 2021).
In this work, we consider the archetype defect in graphene: vacancies Ugeda et al. (2010, 2011). Vacancies, as well as hydrogenation Banhart et al. (2011); Gao et al. (2018) which induce the same type of lattice defects, are both common and controllable. Graphene vacancies create critical defect states pinned at the zero-energy Dirac point by the chiral symmetry Evers and Mirlin (2008); Gade (1993); Altland and Simons (1999); Fukui (2000); Ostrovsky et al. (2006, 2007); Pereira et al. (2006); Ramires and Lado (2019). However, the disorder-origin of the resulting zero-energy flat band raises a strong initial hesitation about its ability to generate physics beyond the localized disorder regime, including many-body instabilities or transport properties.
We use a space-resolved version of the Fubini-Study quantum metric Marzari and Vanderbilt (1997) and show that vacancies in graphene introduce strong long-range correlations. In flat bands in crystalline systems, such a finite quantum metric has recently been shown to generate superconducting phase coherence Peotta and Törmä (2015); Peotta et al. ; Julku et al. (2016); Kopnin et al. (2011), as well as being needed to assess the existence of fractional excitations Varjas et al. (2022). This includes the plethora of phenomena recently discovered in magic-angle twisted or strongly biased bi- and trilayer graphene due to its narrow crystalline flat bands Cao et al. (2018); Hao et al. (2021); Park et al. (2021); Zhou et al. (2021, 2022); Kim et al. (2022); Andrei and MacDonald (2020). A finite quantum metric has also recently been shown to give a nonlinear anomalous Hall effect Gao et al. (2023); Wang et al. (2023). Our results establish that these discovered phenomena in crystalline systems can also be accessible by simply using localized disorder.
Space-resolved quantum metric.— In crystalline systems, due to translational invariance, Bloch’s states $u_{k}^{\alpha}$ form an eigenbasis for the Hamiltonian, with $\alpha$ labeling bands and $k$ the momentum of the quantum state. For a given filling fraction, the quantum metric then reads Peotta and Törmä (2015); Marzari and Vanderbilt (1997):
@@
With disorder present, a real space formulation of the quantum metric is instead required. The origin of such can be traced back to work defining maximally localized Wannier functions Marzari and Vanderbilt (1997); Resta and Sorella (1999); Peotta et al. ; Tam and Peotta . From there the quantum metric can be identified as the reverse Fourier transform of Eq. (1) Marzari and Vanderbilt (1997):
@@
with multiplications by the position operator $\mathbf{r}$ replacing momenta derivatives and $P$ being the projector on the occupied states. Due to the trace, $\Omega_{\mu\nu}$ can be computed over any basis of the Hilbert space. In crystalline systems, the usual basis is the Bloch’s states, as in Eq. (1). In non-crystalline systems we find it useful to compute the trace in the site (atomic) basis, $|i\rangle$, to be able to define a space-resolved quantum metric marker: $\langle i|P\mathbf{r}_{\mu}(1-P)\mathbf{r}_{\nu}P|i\rangle$. Together with the recently developed local Chern marker Bianco and Resta (2011); Marsal et al. (2020), this provide a fully space-resolved notion of the quantum geometric tensor Resta (2002).
@@
Graphene vacancies.—With a space-resolved quantum metric now at hand we apply it to show how systems with localized disorder can host a non-trivial quantum metric. In particular, we focus on a random distribution of vacancies in a graphene monolayer. We model graphene using the usual tight-binding Hamiltonian on honeycomb lattice, encoding the conducting $p_{z}$ electron on each site interacting through nearest-neighbor hopping and with unit cell length $a$ Castro Neto et al. (2009). We then introduce vacancies by removing a number $N_{vac}$ of sites $(v_{i})$ from the lattice, setting all hopping terms to the vacancy site to zero, resulting in the Hamiltonian
@@
Finally, we remove the trivially isolated orbitals $c_{v_{i}}$ from the Hilbert space. We typically consider large systems with open boundary conditions to avoid periodic images, using up to $N=5400$ sites and check our key results for spurious finite size effects. As a vacancy defect only acts on bonds (by removing them), it preserves the chiral symmetry of graphene and is thus a type of chiral disorder, which has been studied both generically and in the honeycomb lattice Evers and Mirlin (2008); Gade (1993); Altland and Simons (1999); Fukui (2000); Ostrovsky et al. (2006, 2007). Each vacancy removes one state from the spectrum, thereby leaving one state that has to be its own chiral partner, thus pinned to zero energy. This defect state is localized on the opposite sublattice to that of the vacancy (see Fig. 1a) and, often missed, it is multifractal, meaning its wave function amplitude decreases as a power-law with distance from the vacancy, in contrast to both fully localized defect states and extended Bloch states. While carbon vacancies may also induce lattice reconstruction Meyer et al. (2008), hydrogenation also effectively remove one $p_{z}$ electron and occur primarily without lattice reconstruction and is experimentally controllable  Banhart et al. (2011); Gao et al. (2018). Technically, adatoms create an onsite perturbation, breaking chiral symmetry, but hydrogen absorbs strongly and still closely resembles a vacancy Casolo et al. (2009); González-Herrero et al. (2016) and can thus be modeled by Eq. (3).
@@
We can further understand our results by noting that chiral disorder (disorder preserving chiral symmetry) is known to create multifractal states, meaning that they are neither exponentially localized nor fully delocalized Gade (1993); Altland and Simons (1999); Fukui (2000); Ostrovsky et al. (2006, 2007); Häfner et al. (2014). This implies a close relationship with the critical regime of the Anderson metal-insulator localization transition Evers and Mirlin (2008). Indeed, vacancies in graphene, irrespective of the disorder concentration, can be expected to behave as a multifractal insulator Cuevas and Kravtsov (2007), with enhanced correlations between states with small energy separation. This is consistent with the enhancement of $\Omega_{\alpha}$ at small energies in Fig. 2(b). To deepen the analogy with the Anderson localization transition we write, for a pair of chiral symmetric states $\alpha$ and $\overline{\alpha}$ with energy splitting $\omega$:
@@
$\Omega_{\alpha}$ is now a two-point and two-state correlator whose exact value depends on the exact position of the vacancies. By utilizing a variational approach developed for Anderson localization Houghton et al. (1980), we find that, on average, the strongest contribution to the integral in Eq. (8) comes from pairs of vacancies and only depends on the energy splitting $\omega$ such that $\Omega_{\alpha}\propto\log^{d+1}\left(\frac{\delta}{\omega}\right)$, where $d=2$ is the system dimension. This variational approach is valid for small energy separation $\omega\ll\delta$ Houghton et al. (1980); Feigel’man et al. (2010), which is the case for a vacancy band. In Fig. 2(b) we also include a fit to the $\log^{p}$ functional form, finding $p=d+1\approx 2.6$, which gives good agreement. For this fit we exclude vacancies with $\Omega_{\alpha}>\frac{\sqrt{3}}{4}Na^{2}$, the area of the sample, which sit close to the sample edge and thus result in strong fluctuations in $\Omega_{\alpha}$ due to finite size effects. These arguments establish that the strong quantum metric enhancement we observe is a consequence of the multifractality of hybridized vacancy states. Notably, the quantum metric enhancement is as strongest for vacancies far apart, a seemingly non-intuitive consequence of the multifractalness.
@@
In Fig. 3(a) we plot the enhancement of the quantum metric (beyond the pristine bulk value $\Omega_{0}$) at half-filling as a function of the vacancy concentration $n^{*}=N_{vac}/N$. The quantum metric (blue points) initially grows with defect concentration, but then saturates and eventually decreases. To understand the small to moderate concentration regime, we imagine a situation where all vacancies are separated by the average distance $d^{*}=1/\sqrt{n^{*}}$. An increase in concentration then results in more pairs contributing to the quantum metric, but with smaller contributions each, with the quantum metric enhancement given by $\Omega^{*}\propto\sum_{\mathrm{pairs\ i,j}}(\mathbf{r}_{i}-\mathbf{r}_{j})^{2}=\frac{Sn^{*}}{2}\times d^{*2}=\frac{S}{2}$, with $S$ the surface area of the sample times a constant. We plot $\Omega^{*}$ in Fig. 3(a) (dashed horizontal line) and note that the quantum metric is always substantially larger than this averaged value. Hence we conclude that “rare” long-distance pairs must be responsible for the quantum metric increase with concentration. At a concentration $\sim 5\%$, the effect of these “rare” pairs saturates and they are instead superseded by closer neighboring pairs, eventually leading to neighboring vacancies, which explains the diminishing quantum metric at large concentrations.
@@
In summary, by introducing a space-resolved formulation of the quantum metric we show that random vacancies in a graphene monolayer generates strong long-range correlations at half filling, primarily from pairs of chiral symmetric vacancy states situated on different sublattices and with a small energy splitting, or equivalently, large spatial separation. Such “rare” long-distance pairs provide an anomalous contribution, particularly at finite concentrations. By utilizing similarities with the Anderson localization transition, we show that the multifractal vacancy wave functions is a key component of the quantum metric enhancement, but we also find that chiral symmetry is an additional important ingredient. A large quantum metric establishes the highly nontrivial nature of the flat vacancy band in graphene. Possible extensions include studying the space-resolved quantum metric for other non-crystalline systems hosting critical states, including quasicrystals, fractals Iliasov et al. , or amorphous systems with vacancies. The spatial repartition of the quantum metric may also provide supplementary information about the electronic properties of the system. We also note that some similar phenomenology was very recently also demonstrated in a Kitaev spin liquid on a honeycomb lattice with vacancies Kao et al. .
Acknowledgements.— We thank T. Löthman, R. Arouca, L. Baldo, and A. Bhattacharya for interesting and fruitful discussions related to this work. We acknowledge funding from the Knut and Alice Wallenberg Foundation through the Wallenberg Academy Fellows program and the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC-2022-CoG, Grant agreement No. 101087096). Our calculations were performed using the Python package kwant Groth et al. (2014) and our plots using matplotlib Hunter (2007). The code used for the numerical calculations and the data shown in the manuscript is available at Ref. Marsal and Black-Schaffer (2023).

### v2

@@ Abstract
The two-dimensionality and linear Dirac dispersion of graphene generate remarkable physical properties novoselov2004 ; novoselov2005 ; castro2009 ; wehling2014 . However, the vanishing density of states at half-filling, at the Dirac point, causes a natural insensitivity to electron interactions and its generated many-body instabilities, otherwise key concepts in condensed matter. But, it makes graphene sensitive to the presence of structural defects castro2009 ; banhart2011 ; yang2018 . Defects often create localized states that affects the conductivity, including in graphene stauber2007 ; wehling2010 , and can even turn a metal into an insulator evers2008 , while defect states that are not fully localized have been shown to favor some long-range correlations feigelman2007 ; burmistrov2012 ; burmistrov2021 .
In this work, we consider the archetype defect in graphene: vacancies ugeda2010 ; ugeda2011 . Vacancies, as well as hydrogenation banhart2011 ; yang2018 which induce the same type of lattice defects, are both common and controllable. Graphene vacancies create critical defect states pinned at the zero-energy Dirac point by the chiral symmetry evers2008 ; gade1993 ; altland1999 ; fukui2000 ; ostrovsky2006 ; ostrovsky2007 ; pereira2006 ; ramires1019 . However, the disorder-origin of the resulting zero-energy flat band raises a strong initial hesitation about its ability to generate physics beyond the localized disorder regime, including many-body instabilities or transport properties.
We use a space-resolved version of the Fubini-Study quantum metric marzari1997 and show that vacancies in graphene introduce strong long-range correlations. In flat bands in crystalline systems, such a finite quantum metric has recently been shown to generate superconducting phase coherence peotta2015 ; peotta2023 ; julku2016 ; Kopnin2011 , as well as being needed to assess the existence of fractional excitations varjas2022 . This includes the plethora of phenomena recently discovered in magic-angle twisted or strongly biased bi- and trilayer graphene due to its narrow crystalline flat bands Cao2018SC ; Hao2021 ; Park2021 ; Zhou2021SC ; zhou2022 ; kim2022 ; andrei2020 . A finite quantum metric has also recently been shown to give a nonlinear anomalous Hall effect Gao2023 ; Wang2023 . Our results establish that these discovered phenomena in crystalline systems can also be accessible by simply using localized disorder.
Space-resolved quantum metric.— In crystalline systems, due to translational invariance, Bloch’s states $u_{k}^{\alpha}$ form an eigenbasis for the Hamiltonian, with $\alpha$ labeling bands and $k$ the momentum of the quantum state. For a given filling fraction, the quantum metric then reads peotta2015 ; marzari1997 :
@@
With disorder present, a real space formulation of the quantum metric is instead required. The origin of such can be traced back to work defining maximally localized Wannier functions marzari1997 ; resta1999 ; peotta2023 ; tam2023 . From there the quantum metric can be identified as the reverse Fourier transform of Eq. (1) marzari1997 :
@@
with multiplications by the position operator $\mathbf{r}$ replacing momenta derivatives and $P$ being the projector on the occupied states. Due to the trace, $\Omega_{\mu\nu}$ can be computed over any basis of the Hilbert space. In crystalline systems, the usual basis is the Bloch’s states, as in Eq. (1). In non-crystalline systems we find it useful to compute the trace in the site (atomic) basis, $|i\rangle$, to be able to define a space-resolved quantum metric marker: $\langle i|P\mathbf{r}_{\mu}(1-P)\mathbf{r}_{\nu}P|i\rangle$. Together with the recently developed local Chern marker bianco2011 ; marsal2020 , this provide a fully space-resolved notion of the quantum geometric tensor resta2002 .
@@
Graphene vacancies.—With a space-resolved quantum metric now at hand we apply it to show how systems with localized disorder can host a non-trivial quantum metric. In particular, we focus on a random distribution of vacancies in a graphene monolayer. We model graphene using the usual tight-binding Hamiltonian on honeycomb lattice, encoding the conducting $p_{z}$ electron on each site interacting through nearest-neighbor hopping and with unit cell length $a$ castro2009 . We then introduce vacancies by removing a number $N_{vac}$ of sites $(v_{i})$ from the lattice, setting all hopping terms to the vacancy site to zero, resulting in the Hamiltonian
@@
Finally, we remove the trivially isolated orbitals $c_{v_{i}}$ from the Hilbert space. We typically consider large systems with open boundary conditions to avoid periodic images, using up to $N=5400$ sites and check our key results for spurious finite size effects. As a vacancy defect only acts on bonds (by removing them), it preserves the chiral symmetry of graphene and is thus a type of chiral disorder, which has been studied both generically and in the honeycomb lattice evers2008 ; gade1993 ; altland1999 ; fukui2000 ; ostrovsky2006 ; ostrovsky2007 . Each vacancy removes one state from the spectrum, thereby leaving one state that has to be its own chiral partner, thus pinned to zero energy. This defect state is localized on the opposite sublattice to that of the vacancy (see Fig. 1a) and, often missed, it is multifractal, meaning its wave function amplitude decreases as a power-law with distance from the vacancy, in contrast to both fully localized defect states and extended Bloch states. While carbon vacancies may also induce lattice reconstruction meyer2008 , hydrogenation also effectively remove one $p_{z}$ electron and occur primarily without lattice reconstruction and is experimentally controllable  banhart2011 ; yang2018 . Technically, adatoms create an onsite perturbation, breaking chiral symmetry, but hydrogen absorbs strongly and still closely resembles a vacancy casolo2009 ; gonzales2016 and can thus be modeled by Eq. (3).
@@
We can further understand our results by noting that chiral disorder (disorder preserving chiral symmetry) is known to create multifractal states, meaning that they are neither exponentially localized nor fully delocalized gade1993 ; altland1999 ; fukui2000 ; ostrovsky2006 ; ostrovsky2007 ; hafner2014 . This implies a close relationship with the critical regime of the Anderson metal-insulator localization transition evers2008 . Indeed, vacancies in graphene, irrespective of the disorder concentration, can be expected to behave as a multifractal insulator cuevas2007 , with enhanced correlations between states with small energy separation. This is consistent with the enhancement of $\Omega_{\alpha}$ at small energies in Fig. 2(b). To deepen the analogy with the Anderson localization transition we write, for a pair of chiral symmetric states $\alpha$ and $\overline{\alpha}$ with energy splitting $\omega$:
@@
$\Omega_{\alpha}$ is now a two-point and two-state correlator whose exact value depends on the exact position of the vacancies. By utilizing a variational approach developed for Anderson localization houghton1980 , we find that, on average, the strongest contribution to the integral in Eq. (8) comes from pairs of vacancies and only depends on the energy splitting $\omega$ such that $\Omega_{\alpha}\propto\log^{d+1}\left(\frac{\delta}{\omega}\right)$, where $d=2$ is the system dimension. This variational approach is valid for small energy separation $\omega\ll\delta$ houghton1980 ; feigelman2010 , which is the case for a vacancy band. In Fig. 2(b) we also include a fit to the $\log^{p}$ functional form, finding $p=d+1\approx 2.6$, which gives good agreement. For this fit we exclude vacancies with $\Omega_{\alpha}>\frac{\sqrt{3}}{4}Na^{2}$, the area of the sample, which sit close to the sample edge and thus result in strong fluctuations in $\Omega_{\alpha}$ due to finite size effects. These arguments establish that the strong quantum metric enhancement we observe is a consequence of the multifractality of hybridized vacancy states. Notably, the quantum metric enhancement is as strongest for vacancies far apart, a seemingly non-intuitive consequence of the multifractalness.
@@
In Fig. 3(a) we plot the enhancement of the quantum metric (beyond the pristine bulk value $\Omega_{0}$) at half-filling as a function of the vacancy concentration $n^{*}=N_{vac}/N$. The quantum metric (blue points) initially grows with defect concentration, but then saturates and eventually decreases. To understand the small to moderate concentration regime, we imagine a situation where all vacancies are separated by the average distance $d^{*}=1/\sqrt{n^{*}}$. An increase in concentration then results in more pairs contributing to the quantum metric, but with smaller contributions each, with the quantum metric enhancement given by $\Omega^{*}\propto\sum_{\mathrm{pairs\ i,j}}(\mathbf{r}_{i}-\mathbf{r}_{j})^{2}=\frac{Sn^{*}}{2}\times d^{*2}=\frac{S}{2}$, with $S$ the surface area of the sample times a constant. We plot $\Omega^{*}$ in Fig. 3(a) (dashed horizontal line) and note that the quantum metric is always substantially larger than this averaged value. Hence we conclude that “rare” long-distance pairs must be responsible for the quantum metric increase with concentration. A similar phenomenology has been predicted in flat bands made of compact localized states bouzerar2022 . At a concentration $\sim 5\%$, the effect of these “rare” pairs saturates and they are instead superseded by closer neighboring pairs, eventually leading to neighboring vacancies, which explains the diminishing quantum metric at large concentrations.
@@
In summary, by introducing a space-resolved formulation of the quantum metric we show that random vacancies in a graphene monolayer generates strong long-range correlations at half filling, primarily from pairs of chiral symmetric vacancy states situated on different sublattices and with a small energy splitting, or equivalently, large spatial separation. Such “rare” long-distance pairs provide an anomalous contribution, particularly at finite concentrations. By utilizing similarities with the Anderson localization transition, we show that the multifractal vacancy wave functions is a key component of the quantum metric enhancement, but we also find that chiral symmetry is an additional important ingredient. A large quantum metric establishes the highly nontrivial nature of the flat vacancy band in graphene. Possible extensions include studying the space-resolved quantum metric for other non-crystalline systems hosting critical states, including quasicrystals, fractals iliasov2023 , or amorphous systems with vacancies. The spatial repartition of the quantum metric may also provide supplementary information about the electronic properties of the system. We also note that a recent work has discussed the relation between the quantum metric and the Berry curvature in the Haldane model in the presence of vacancies romeral2024 and that some similar phenomenology was very recently also demonstrated in a Kitaev spin liquid kao2023 .
Acknowledgements.— We thank T. Löthman, R. Arouca, L. Baldo, and A. Bhattacharya for interesting and fruitful discussions related to this work. We acknowledge funding from the Knut and Alice Wallenberg Foundation through the Wallenberg Academy Fellows program and the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC-2022-CoG, Grant agreement No. 101087096). Our calculations were performed using the Python package kwant groth2014 and our plots using matplotlib hunter2007 . The code used for the numerical calculations and the data shown in the manuscript is available at Ref. marsal2023 .